# Building Bridges, Not Walls: Advancing Interpretability by Unifying Feature, Data, and Model Component Attribution

### v2

@@ 7 Alternative Views
While we propose a unified view of attribution methods, alternative views exist for characterizing them.
One alternative view suggests that feature, data, and component attribution represent fundamentally distinct research directions, each with its own terminology and research community. Despite using similar methods, these fields differ in their motivations and methodological approaches. The term “attribution” carries unique meanings in each case. Feature attribution has historically been synonymous with “explainable AI,” data attribution was part of “data-centric AI” focusing on concrete applications like data selection, and component attribution has emerged recently as one branch of “mechanistic interpretability” to reverse-engineer algorithms in foundation models. This rather independent evolution has produced distinct best practices within each domain (Guidotti et al., 2018; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024) and separate communities that may seem to resist unification. Nevertheless, we argue that meaningful unification across these three aspects is possible and will help advance interpretability research.
@@
Researchers have also proposed various taxonomies within each type of attribution. Feature attribution methods, for example, were classified as model-agnostic or model-specific approaches (Arrieta et al., 2020). Model-agnostic methods treat models as black boxes and generate attributions based solely on input-output relationships, while model-specific methods rely on known model architectures. Although these categorizations aid method selection, they have limited generalizability to other aspects of attribution. Component attribution methods, for instance, are inherently model-specific since they analyze defined model components. Our unification approach extends beyond attribution-specific frameworks to take a broader, cross-aspect perspective. This comprehensive view uncovers connections that remain obscured when examining methods in isolation. Consider gradient-based approaches: while methods from all three types of attribution employ gradients, they compute them for different purposes with different objectives. Thus, though aspect-specific categorizations offer valuable insights, our cross-aspect framework provides a holistic understanding of the attribution landscape.
@@
The recent work by Saphra and Wiegreffe (2024) offers another alternative view on attribution methods by examining the evolving definitions of mechanistic interpretability, a parent field of component attribution. Their study highlights how different communities within interpretability research assign distinct meanings to attributions, often guided by historical, cultural, and methodological differences, underscoring a semantic drift in how terminologies are understood. This perspective aligns with our unification position in that it recognizes attribution methods as a function of community-specific definitions and epistemic commitments, rather than strictly independent methodologies.
@@ 8 Conclusion
In this paper, we have presented a unified view on three traditionally separate attribution methods for interpretability: feature attribution, data attribution, and component attribution. While these methods have evolved independently to explain different aspects of AI systems we demonstrate that they share fundamental technical building blocks such as perturbations, gradients, and linear approximations. This unified view not only bridges the fragmented landscape of attribution methods but also identifies common challenges and opportunities for cross-attribution knowledge transfer, ultimately working toward more comprehensive approaches to advancing AI interpretability research.
9 Acknowledgements
@@ Appendix B Summary of Methods
As we discussed in the main body, attribution methods across features, data, and components can be categorized into three main approaches: perturbation-based methods, gradient-based methods, and linear approximation methods. Perturbation-based methods measure how a model’s output changes when modifying specific elements, whether they are input features, training data points, or model components. To capture interactions between multiple elements, all three types of attribution methods employ common mathematical tools, such as the Shapley value from game theory. Gradient-based methods analyze model behavior by leveraging gradients to provide insights into the model’s sensitivity to small input changes. Gradients bridge model behavior and the elements we wish to attribute without perturbations. Attributions are achieved in different types of gradients: computing gradients of model outputs with respect to input features to quantify feature importance, calculating gradients of loss functions with respect to specific training data points to analyze data influence, or using gradients to approximate the effects of modifying model components. Linear approximation methods fit linear models to approximate complex model behaviors. The inputs to these linear models can be input features, training data points, or model components. In some cases, binary indicators replace the actual elements as inputs to simplify the approximation.
In total, thousands of attribution methods of all three types have been proposed making a comprehensive literature summary infeasible. In Table 1, we summarize the attribution methods discussed in this paper, which we believe are the representative ones and align with the unified view we presented. Most of the empty cells in Table 1 (labeled as “–”) represent methods that we believe are promising but have not yet been explored in the literature, as discussed in § 6.3. These represent research ideas that have been verified in one attribution type but remain unexplored in others. The exception is mask learning for data attribution, which we consider less promising because learning a high-dimensional mask of size jointly with the model would be infeasible when the model has not been trained.
@@ C.4 Gradient-Based Feature Attribution
Gradients for feature attribution are very different from those used in model training. For a model with parameters , gradients of the loss function are taken with respect to the parameters () to guide parameter updates during training. In feature attribution, gradients of the model’s output are taken with respect to input features () to quantify each feature’s contribution to the model’s output. Gradient-based methods have been widely adopted for feature attribution because of their computational efficiency. They typically require only a single forward and backward pass through the model to compute the gradients, and require no additional perturbation or linear model fitting, which makes them particularly suitable for real-time applications and large-scale datasets. On the other hand, gradient-based methods have two key limitations. They require access to model parameters and only work with differentiable models. Additionally, the gradient results can be nonrobust as we discussed in Section 6.2. Despite these challenges, gradient-based methods remain a fundamental tool in the feature attribution toolkit.
@@ Appendix D Detailed Discussion of Data Attribution Methods
Unlike feature attribution, data attribution methods trace back to the training phase and quantifies the training data’s influence on the model’s output, but they similarly fall into three categories. Perturbation-based methods assess training data importance by observing changes in model behavior when training samples are removed or modified. These methods provide accurate results but can be computationally expensive as they require retraining the model multiple times. Complete retraining-based methods like LOO are often used as a ground truth for evaluating other data attribution methods. Gradient-based methods utilize the model’s gradients evaluated at the training data points and the test data point to quantify the influence of the training data points on the test data point. These methods avoid the computational cost of retraining the model but may face the challenges like the non-convexity of the loss landscape or the difficulty in computing the Hessian matrix efficiently. Linear approximation methods construct interpretable models that approximate how training data affects model behavior. The linear model operates on the entire training dataset, which can be suprisingly accurate but also heavy to train. Methods from different categories have their own strengths, limitations, and use cases. We now extend the discussion of some prominent methods mentioned in the main text in more detail and provide discussions of some additional methods.
@@ D.4 Tracing Training Path for Data Attribution
Tracing training path for data attribution provides valuable insights of training dynamics while avoiding limitations of LOO and IF. Besides TracIn, there are other methods that trace training dynamics that provide more accurate attribution results but they are all more computationally expensive than those considering only final model parameters like IF.
@@ Appendix E Detailed Discussion of Component Attribution Methods
Unlike feature and data attribution, component attribution methods analyze the internal mechanisms of models by attributing model behavior to specific architectural components like neurons, layers, or attention heads. These methods similarly fall into three categories. Perturbation-based methods assess component importance by observing changes in model behavior when specific components are modified, resulting in various forms of causal mediation analysis. Gradient-based methods utilize gradients with respect to component activations to approximate the component importance in causal mediation analysis. Linear approximation methods construct linear models that directly approximate how components affect model behavior. Methods from different categories have their own strengths and limitations. We now extend the discussion of some prominent methods mentioned in the main text in more detail and provide discussions of some additional methods.
@@ E.1 Perturbation-Based Component Attribution
The causal mediation analysis is frequently referred to as activation patching, wherein activations of the specific component from the clean run are patched into the corrupted run to ascertain if those activations are sufficient and necessary to retrieve the desired output. Activation perturbations can consist of zero ablations [Olsson et al., 2022, Geva et al., 2023], mean ablations [Wang et al., 2022], smoothed Gaussian noising Meng et al. [2022], interchange interventions [Geiger et al., 2021], learned ablations [Li and Janson, 2024]. In all cases, the dataset used to generate the activations must be chosen to elicit the desired model behavior, with a matching metric that measures the success of the behavior.
@@ E.3 Generalizing the Definitions of Components
Appendix F Challenges in Attribution Methods
F.1 Computation Challenges of Attribution Methods
@@
Computation challenges present substantial barriers that often prevent attribution methods from being applied to large-scale AI models, such as the foundation models with billions of parameters. For perturbation-based methods, the curse of dimensionality makes a comprehensive analysis intractable when the number of required perturbations is large. For example, the full power set perturbation. This holds for all three types of attributions including high-dimensional inputs, large training datasets, and models with numerous components. Game-theoretic methods face particular difficulties, as exact computation of the Shapley value is often prohibitively expensive and requires approximation techniques like Monte Carlo sampling. The computational burden is also severer for data attribution methods, which require model retraining for each perturbation. Gradient-based methods are more practical for large-scale models. However, gradients essentially only provide first-order approximations of model behavior, which are inadequate to capture complex model behaviors and more sophisticated gradient formulations are needed for better attribution results. For instance, TracIn require aggregating gradients across multiple stages, while IF demand computation of second-order Hessian matrices, leading to increased computational overhead. Linear approximation methods also face computational hurdles in achieving high-quality approximations. Model behavior can be complex, requiring numerous data points and model evaluations to establish sufficient data for learning accurate linear models. Furthermore, for all three types, most attribution methods must compute results separately for each new test data point, creating additional computational strain when attribution analysis is needed for large datasets.
F.2 Consistency of Attribution Methods
@@
The consistency problem in attribution methods is a significant concern. This challenge is also prevalent across due to variability introduced in sampling, learning processes with stochastic optimization, and also non-trivial hyperparameters. When attribution involves sampling, such as the Monte Carlo sampling in some perturbation-based methods to avoid the full power set perturbation, the inherent randomness leads to varying attribution results. Besides, when attribution involves learning processes with stochastic optimization, as seen in mask-learning perturbation and linear approximation methods, different learning outcomes yield inconsistent attribution results. Many attribution methods rely on hyperparameters that can lead to different attribution outcomes. These include sampling parameters, such as the number of samples used for computing Shapley values. They also include optimization hyperparameters for various learning approaches, such as learning rates and number of steps in linear approximation methods. Additionally, approximation hyperparameters are needed for quantities that are computationally challenging to calculate directly, such as dampening factors for inverse Hessian-vector products. Further variability is introduced through fundamental design choices, such as the selection of perturbation type in perturbation-based methods, where options include mean perturbation, zero perturbation, and random perturbation. While these different approaches should theoretically produce similar results based on their underlying principles, in practice they often yield notably different attributions.
F.3 Evaluation of Attribution Methods

### v3

@@ 5 Alternative views
Several alternative views exist that (partially) oppose the unified perspective presented in this paper. We discuss their merits and connections to our work below.
One alternative view holds that different attributions represent distinct research areas that have benefited from independent evolution in the past, and although they share some ideas, they will continue to benefit from developing separately. FA has historically been synonymous with ”explainable AI,” DA has been part of ”data-centric AI” focusing on applications such as data selection, and CA has emerged as a branch of ”mechanistic interpretability” aimed at reverse-engineering algorithms in foundation models. This independent evolution has produced distinct motivations and best practices within each area (Guidotti et al., 2018; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024). While this view does not strongly oppose unification, attribution methods can indeed be more easily compared with other works from their respective fields when viewed separately. For example, FA can be compared with inherently interpretable models (Koh et al., 2020) for feature importance, DA can be compared with data coresets (Mirzasoleiman et al., 2020) for data selection, and CA can be compared with model pruning (Sun et al., 2023). Nevertheless, we demonstrate in the second part of our position that meaningful unification across attributions is possible and can inspire new research directions. We emphasize the unified perspective and its benefits throughout this paper.
@@
There are also alternative views on the taxonomy of attribution methods. First, researchers have proposed other taxonomies within each type of attribution. For example, FA methods have been classified as either model-agnostic or model-specific (Arrieta et al., 2020). Although these categorizations aid method selection, they are not applicable across all attribution types. CA methods, for instance, are inherently model-specific since they analyze defined model components. Second, one might argue that our three-way categorization of FA, DA, and CA into perturbation-based, gradient-based, and linear-approximation methods does not apply to all attribution methods. While this is true – there are surrogate models for approximation beyond linear (Sokol and Flach, 2020) for FA (Section C.5), simulation-based (Chai et al., 2024) and reinforcement learning-based (Yoon et al., 2020) methods for DA (Section D.6), and sparse autoencoder-based methods for CA (Bricken et al., 2023) (Section E.3) – our categorization captures the majority of representative methods. In summary, our unification extends beyond attribution-specific categorizations and highlights broader, cross-attribution similarities. It is relatively comprehensive and uncovers connections that remain obscured when examining attributions in isolation. For example, while methods from all three attribution types use gradients, they compute them for different purposes with different objectives.
@@
Recent work by Saphra and Wiegreffe (2024) offers seemingly another alternative view on attributions, highlighting that different communities assign distinct meanings to attributions, often guided by historical, cultural, and methodological differences. However, by recognizing that attributions carry community-specific definitions with epistemic commitments, rather than being strictly independent methodologies, this view is actually aligned with our unification position.
@@ 6 Conclusion
In this paper, we present a unified view of attribution methods from explainable AI (e.g., FA), data-centric AI (e.g., DA), and mechanistic interpretability (e.g., CA). We demonstrate that, while these attributions evolved largely independently to explain different aspects of AI systems, they differ mainly in perspective (i.e., examining input features, training data, or model components) and share core techniques (i.e., perturbations, gradients, and linear approximations). This unified view not only bridges the fragmented landscape of attribution methods but also inspires new directions in interpretability research through addressing common challenges and cross-attribution innovation and in broader AI research through model editing, steering, and regulation.
7 Acknowledgments
@@ Appendix B Summary of methods
As we discussed in the main paper, methods for feature, data, and component attribution can be categorized based on their technique: perturbation-based methods, gradient-based methods, and linear approximation methods. Perturbation-based methods measure how a model’s output changes when modifying specific elements, whether they are input features, training data points, or model components. To capture interactions between multiple elements, all three types of attributions employ common mathematical tools, such as the Shapley value from game theory. Gradient-based methods analyze model behavior by leveraging gradients to provide insights into the model’s sensitivity to small input changes. Gradients bridge model behavior and the elements we wish to attribute without perturbations. Attributions are generated using different kinds of gradients: computing gradients of model outputs with respect to input features to quantify feature importance, calculating gradients of loss functions with respect to specific training data points to analyze data influence, or using gradients to approximate the effects of modifying model components. Linear approximation methods fit linear models to approximate complex model behaviors. The inputs to these linear models can be input features, training data points, or model components. In some cases, binary indicators replace the actual elements as inputs to simplify the approximation.
In Table 1, we summarize the attribution methods discussed in this paper, which we believe are the representative ones and align with the unified view we presented. Most of the empty cells in Table 1 (labeled as “–”) represent methods that may be promising but have not yet been explored in the literature, as discussed in § 4.1.2. These represent research ideas that have been verified in one attribution type but remain unexplored in others. The exception is mask learning for data attribution, which may be less promising because learning a high-dimensional mask of size jointly with the model would be infeasible when the model has not been trained.
@@ C.3 Perturbation mask learning for feature attribution
Masking model
[Dabkowski and Gal, 2017] pioneered the masking approach for image classification by introducing a masking model that generates pixel masks, aiming to identify a minimal set of features that sufficiently maintain the prediction of the original input. This approach employs a U-Net [Ronneberger et al., 2015] as the masking model, which first encodes the input image and then upsamples it to produce a mask. The masking model acts as the attribution function , where mask values represent feature attribution scores. While initial training is required, the masking model generates masks through a single forward pass at test time, which significantly improves runtime compared to earlier perturbation methods. For learning the mask model, the main challenge is balancing feature minimality and predictive power.
@@ C.4 Gradient-based feature attribution
Gradients for feature attribution are very different from those used in model training. For a model with parameters , gradients of the loss function are taken with respect to the parameters () to guide parameter updates during training. In feature attribution, gradients of the model’s output are taken with respect to input features () to quantify each feature’s contribution to the model’s output. Gradient-based methods have been widely adopted for feature attribution because of their computational efficiency. They typically require only a single forward and backward pass through the model to compute the gradients, and require no additional perturbation or linear model fitting, which makes them particularly suitable for real-time applications and large-scale datasets. On the other hand, gradient-based methods have two key limitations. They require access to model parameters and only work with differentiable models. Additionally, the gradient results can be nonrobust as we discussed in Section 4.1.1. Despite these challenges, gradient-based methods remain a fundamental tool in the feature attribution toolkit.
@@ Appendix D Detailed discussion of data attribution methods
Unlike feature attribution methods, data attribution methods examine the training phase and quantify the training data’s influence on the model’s output. However, data attribution methods can be categorized into the same three categories. Perturbation-based methods assess training data importance by observing changes in model behavior when training samples are removed or modified. These methods provide accurate results but can be computationally expensive as they require retraining the model multiple times. Complete retraining-based methods like LOO are often used as a ground truth for evaluating other data attribution methods. Gradient-based methods utilize the model’s gradients evaluated at the training data points and the test data point to quantify the influence of the training data points on the test data point. These methods avoid the computational cost of retraining the model but may face the challenges like the non-convexity of the loss landscape or the difficulty in computing the Hessian matrix efficiently. Linear approximation methods construct interpretable models that approximate how training data affects model behavior. The linear model operates on the entire training dataset, which can be suprisingly accurate but also heavy to train. Methods from different categories have their own strengths, limitations, and use cases. We now extend the discussion of some prominent methods mentioned in the main text and discuss some additional methods.
@@ D.4 Tracing training path for data attribution
Tracing training path for data attribution provides valuable insights of training dynamics while avoiding limitations of LOO and IF. These methods that trace training dynamics that provide more accurate attribution results but they are all more computationally expensive than those considering only final model parameters like IF.
TracIn
[Pruthi et al., 2020] traces the data influence throughout the entire training process. The method attributes influence by computing dot products between training and test data gradients at each training step from the initial model parameters to the final model parameters at the end of training, accumulating these to capture a training point’s total influence across the training path. This path tracing approach provides valuable insights into training dynamics while avoiding limitations of LOO and IF, such as assigning identical attribution scores to duplicate training data points. TracIn also offers greater flexibility than IF by eliminating the convexity assumption and Hessian matrix computations. On the other hand, its tracing requires storing intermediate model checkpoints during training, increasing both memory usage and computational costs.
@@
D.5 Linear approximation and Datamodel
Datamodel
[Ilyas et al., 2022] applies linear approximation to DA, similar to LIME in FA. It constructs a linear model with coefficients and vectors as inputs, where each input represents a subset of training data. is learned to map any counterfactual subset of training data to output , where is trained on this subset with the given model architecture and training algorithm. The coefficients of thus represent the attribution scores of the training data points. The method’s counterfactual nature enables evaluation of other attribution methods via the Linear Datamodeling Score (LDS), which compares their attribution score rankings to Datamodel’s ranking. While Datamodel can effectively capture model behavior, constructing this large linear model requires extensive counterfactual data obtained by training model on various subsets, making it computationally intensive.
TRAK
[Park et al., 2023] addresses the computational challenges of Datamodel by estimating Datamodels in a transformed space where the learning problem becomes convex and can be approximated efficiently. It further improves efficiency through random projection of model parameters and ensemble attribution results of multiple trained models. Though the ensemble approach still requires some model retraining on different subsets, it achieves high estimation accuracy with significantly fewer retraining iterations than Datamodel. Furthermore, both approaches can be viewed as perturbation-based methods, similar to LIME, as they systematically vary training data to construct linear models.
D.6 Data attribution methods using other techniques
There are data attribution methods that use other techniques beyond the three core techniques we discussed in the main paper to estimate the influence of each training data point on the model’s output.
GPTfluence
[Chai et al., 2024] is a simulation-based method that estimates how individual training examples influence the performance of GPT models. GPTfluence uses a featurized simulation-based approach to model the training trajectory. It begins by collecting training dynamics across multiple runs, including sequences of training batches and performance metrics like loss, BLEU, and ROUGE scores. GPTfluence then trains a simulator that represents both training and test examples using frozen pre-trained embeddings (e.g., from BERT), and learns multiplicative and additive influence factors through inner products between these embeddings. The simulator models performance at each step using an n-th order Markov process that conditions on previous steps, allowing it to predict the full evolution of test metrics over training.
Data Valuation using Reinforcement Learning (DVRL)
[Yoon et al., 2020] uses reinforcement learning to estimate the influence of each training data point on the model’s output. DVRL jointly trains a data value estimator and a task predictor model. At each iteration, the data value estimator assigns selection probabilities to training examples, determining which examples are used to update the predictor model. The quality of these selections is evaluated using a small validation set, and a RL signal based on the validation performance is used to update the data value estimator. This allows DVRL to dynamically learn which samples are most useful for improving the model’s performance.
@@ Appendix E Detailed discussion of component attribution methods
Unlike feature and data attribution methods, component attribution methods analyze the internal mechanisms of models by attributing model behavior to specific architectural components like neurons, layers, or attention heads. These methods similarly fall into three categories. Perturbation-based methods assess component importance by observing changes in model behavior when specific components are modified, resulting in various forms of causal mediation analysis. Gradient-based methods utilize gradients with respect to component activations to approximate the component importance in causal mediation analysis. Linear approximation methods construct linear models that directly approximate how components affect model behavior. Methods from different categories have their own strengths and limitations. We now extend the discussion of some prominent methods mentioned in the main paper and provide discussions of some additional methods.
@@ E.1 Perturbation-based component attribution
Causal mediation analysis is frequently referred to as activation patching, wherein activations of the specific component from the clean run are patched into the corrupted run to ascertain if those activations are sufficient and necessary to retrieve the desired output. Activation perturbations can consist of zero ablations [Olsson et al., 2022, Geva et al., 2023], mean ablations [Wang et al., 2022], smoothed Gaussian noising Meng et al. [2022], interchange interventions [Geiger et al., 2021], learned ablations [Li and Janson, 2024]. In all cases, the dataset used to generate the activations must be chosen to elicit the desired model behavior, with a matching metric that measures the success of the behavior.
@@ E.3 Generalizing the definitions of components
Appendix F Overcoming attribution challenges through a unified lens
F.1 Improving computational efficiency
@@
Computation challenges present substantial barriers that often prevent attribution methods from being applied to large-scale AI models, such as the foundation models with billions of parameters. For perturbation-based methods, the curse of dimensionality makes a comprehensive analysis intractable when the number of required perturbations is large. For example, the full power set perturbation. This holds for all three types of attributions including high-dimensional inputs, large training datasets, and models with numerous components. Game-theoretic methods face particular difficulties, as exact computation of the Shapley value is often prohibitively expensive and requires approximation techniques like Monte Carlo sampling. The computational burden is also severer for data attribution methods, which require model retraining for each perturbation. Gradient-based methods are more practical for large-scale models. However, gradients essentially only provide first-order approximations of model behavior, which are inadequate to capture complex model behaviors and more sophisticated gradient formulations are needed for better attribution results. For instance, TracIn require aggregating gradients across multiple stages, while IF demand computation of second-order Hessian matrices, leading to increased computational overhead. Linear approximation methods also face computational hurdles in achieving high-quality approximations. Model behavior can be complex, requiring numerous data points and model evaluations to establish sufficient data for learning accurate linear models. Furthermore, for all three types, most attribution methods must compute results separately for each new test data point, creating additional computational strain when attribution analysis is needed for large datasets. Therefore, improving computational efficiency calls for better sampling methods to avoid perturbing the entire input space in perturbation-based methods and to learn more accurate linear models for linear approximation methods. Also needed are more efficient gradient computations, especially for higher-order gradients like Hessians. Many works have sought to improve these aspects, with promising examples like Grosse et al. [2023] showing that gradient-based methods can be scaled to large language models (LLMs) with up to 52 billion parameters. However, there is still much room for improvement to make attribution methods more efficient.
F.2 Improving attribution consistency
@@
The consistency problem in attribution methods is a significant concern. This challenge is also prevalent across due to variability introduced in sampling, learning processes with stochastic optimization, and also non-trivial hyperparameters. When attribution involves sampling, such as the Monte Carlo sampling in some perturbation-based methods to avoid the full power set perturbation, the inherent randomness leads to varying attribution results. Besides, when attribution involves learning processes with stochastic optimization, as seen in mask-learning perturbation and linear approximation methods, different learning outcomes yield inconsistent attribution results. Many attribution methods rely on hyperparameters that can lead to different attribution outcomes. These include sampling parameters, such as the number of samples used for computing Shapley values. They also include optimization hyperparameters for various learning approaches, such as learning rates and number of steps in linear approximation methods. Additionally, approximation hyperparameters are needed for quantities that are computationally challenging to calculate directly, such as dampening factors for inverse Hessian-vector products. Further variability is introduced through fundamental design choices, such as the selection of perturbation type in perturbation-based methods, where options include mean perturbation, zero perturbation, and random perturbation. While these different approaches should theoretically produce similar results based on their underlying principles, in practice they often yield notably different attributions. While some gradient-based methods can produce consistent results in a single run when they do not involve sampling or approximations for computationally intensive operations, the consistency among different gradient-based methods varies considerably, which is also a problem for all three techniques and leads to the following evaluation challenges. We see the potential improvement for this challenge including producing more deterministic attribution methods that are less sensitive to random hyperparameters, provide provable guarantees of attribution stability or error bounds. While past works have made some progress in this direction, there were strong assumptions added to the setting, e.g., convexity [Koh and Liang, 2017]. Thus, more efforts are needed to improve consistency in more realistic settings.
F.3 Towards fairer and more practical evaluation